OpenStudy (anonymous):

Evaluate integral of( x sin^{-1}x)/sqrt(1-x^{2})dx

OpenStudy (anonymous):

\[\int\limits \frac{xsin^{-1}x}{sqrt{1-x^{2}}\]

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{xsin^-{1}x}{\sqrt{1-x^2}}dx\]

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{x}{\sqrt{1-x^2}} \cdot \sin^{-1}(x) dx \text{ use integration by parts }\] ---------------------------- \[\int\limits_{}^{}\frac{x}{\sqrt{1-x^2}} dx \text{ } u=1-x^2 => du=-2 x dx => \frac{-1}{2} du=x dx\] \[\int\limits_{}^{}\frac{\frac{-1}{2} du}{\sqrt{u}}=\frac{-1}{2}\int\limits_{}^{} u^\frac{-1}{2} du\] I will let you finish this part ----------------------------- Then ask yourself what is (arcsin(x))' See how this works out for yea Let me know if you need further assistance

OpenStudy (anonymous):

Integrate by parts. \[ u=\sin^{-1} x ,\, dv= \frac x{\sqrt{1-x^2}} \] Work it out

OpenStudy (anonymous):

\[du =\frac 1 {\sqrt{1-x^2}}\\ v= -\sqrt{1-x^2} \] Compute now \[ u v - \int v du \] and you will be done.

OpenStudy (anonymous):

\[x- \sqrt{1-x^2}\sin^-1x\] +c

OpenStudy (anonymous):

thank you sooo much..

OpenStudy (anonymous):

but how did you get\[v=-\sqrt{1-x^2}\]

OpenStudy (freckles):

did you see my post above?

OpenStudy (freckles):

if you follow that you will see

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hld on pls i'm wrkng on it